# On the Importance of Gaussianizing Representations

### v1

@@ 1 Introduction
Our experiments comprehensively demonstrate the general effectiveness of normality normalization in terms of its generalization performance, its strong performance across various common factors of variation such as model width, depth, and training minibatch size, which furthermore serve to highlight why it is effective, its suitability for usage wherever existing normalization layers are conventionally used, and its effect on improving model robustness under random perturbations.
@@ 2.1.1 Overview of the Framework
The normal distribution is at the same time the best possible signal distribution, and the worst possible noise distribution; a result which can be studied in the context of the Gaussian channel (Shannon, 1948), and through the lens of the mutual information game (Cover & Thomas, 2006). In this framework, $X$ and $Z$ denote two independent random variables, representing the input signal and noise, and $Y=X+Z$ is the output. The mutual information between $X$ and $Y$ is denoted by $I\left(X;Y\right)$; $X$ tries to maximize this term, while $Z$ tries to minimize it. Both $X$ and $Z$ can encode their signal using any probability distribution, so that their respective objectives are optimized for.
@@ 2.1.2 Relation to Learning
How might this framework relate to the learning setting? First, previous works have shown that adding noise to the inputs (Bishop, 1995) or to the intermediate activations (Srivastava et al., 2014) of neural networks can be an effective form of regularization, leading to better generalization. Moreover, the mutual information game shows that, among encoding distributions, the normal distribution is maximally robust to random perturbations. Taken together these suggest that encoding activations using the normal distribution is the most effective way of using noise as a regularizer, because a greater degree of regularizing noise in the activations can be tolerated for the same level of corruption.
@@
Finally, there exists a close correspondence between the mutual information between the input and the output of a channel subject to additive Gaussian noise, and the minimum mean-squared error (MMSE) in estimating (or recovering) the input given the output (Guo et al., 2005). This suggests that when Gaussian noise is added to a given layer’s activations, quantifying the attenuation of the noise across the subsequent layers of the network, as measured by the mean-squared error (MSE) relative to the unperturbed activations, provides a direct and measurable proxy for the mutual information between the activations of successive layers in the presence of noise.
@@ 2.3 Maximally Independent Representations
We use these results to motivate the following argument: for a given level of correlation, encouraging normality in the feature representations of units by using normality normalization, would lead to the desirable property of maximal independence between them; in the setting where increased unit-wise normality also lends itself to increased joint normality.
@@ 3 Background: Power Transform
The parameter $\lambda$ is obtained using maximum likelihood estimation (MLE) so that the transformed variable is as normally distributed as possible, by minimizing the following NLL [^2] To simplify the presentation, we momentarily defer the cases $\lambda=0$ and $\lambda=2$, and outline the NLL for $h\geq 0$ only, as the case for $h<0$ follows closely by symmetry. :
@@ 4.2 Additive Gaussian Noise with Scaling
For each input indexed by $i\in\left\{1,\ldots,N\right\}$, during training444 We do not apply additive random noise with scaling at test time. we have $y_{i}=x_{i}+z_{i}\cdot\xi\cdot s$, where $x_{i}$ is the $i$-th input’s post-power transform value, $z_{i}\sim\mathcal{N}\left(0,1\right)$, $\xi\geq 0$ is the noise factor, and $s=\frac{1}{N}\left\lVert\bm{x}-\bar{\bm{x}}\right\rVert_{1}$ represents the zero-centered norm of the post-power transform values, normalized by the sample size $N$.
@@ 5.4 Effectiveness Across Model Configurations
Figure 2 shows that BatchNormalNorm outperforms BatchNorm across varying WideResNet architecture model widths. Of particular note is that BatchNormalNorm shows strong performance even in the regime of relatively small network widths, whereas BatchNorm’s performance deteriorates. This may indicate that for small-width networks, which do not exhibit the Gaussian process limiting approximation attributed to large-width networks (Neal, 1996; Lee et al., 2017; Jacot et al., 2018; Lee et al., 2019), normality normalization provides a correcting effect. This could, for example, be beneficial for hardware-limited deep learning applications.
@@ 5.6 Comparison of Additive Gaussian Noise With Scaling and Gaussian Dropout
The second is additive Gaussian noise, but without scaling by each channel’s minibatch statistics. This corresponds to the proposed method in the case where $s$ is fixed to the mean of a standard half-normal distribution [^8] This value for $s$ precisely mirrors how it is calculated in Algorithm 1, since recall $s=\frac{1}{N}\sum_{i=1}^{N}\lvert x_{i}-\bar{x}\rvert$. , i.e. $s=\sqrt{\frac{2}{\pi}}$ across all channels; and thus does not depend on the channel statistics.
@@ 5.7 Effect of Degree of Gaussianization
Here we consider what effect differing degrees of gaussianization have on model performance, as measured by the proximity of the estimate $\hat{\lambda}$ to its MLE solution, which was given by Equation 5.
We control the proximity to the MLE solution using a parameter $\alpha\in\left[0,1\right]$ in the following equation
@@
where $\alpha=1$ corresponds to the MLE, and decreasing values of $\alpha$ reduce the strength of the gaussianization.
@@ 5.9 Additional Experiments & Analysis
Appendix A.1 & Table 4 demonstrate that models trained using normality normalization are more robust to random noise at test time. This substantiates the applicability of the noise robustness framework presented in Motivation Subsection 2.1, and consequently of the benefit of gaussianizing learned representations.
@@
Appendix A.2 & Figure 10 show that normality normalization increases runtime; with a large deviation at training time than at test time.
@@
Appendix A.3 & Figure 11 show that, at initialization, layer pre-activations are close-to Gaussian regardless of the normalization layer employed; but that only models trained with normality normalization maintain gaussianity throughout training.
@@
Appendix A.4 & Figure 12 demonstrate that normality normalization imbues models with greater joint normality and greater independence between channel features, throughout the layers of a model. This is of value in context of the benefit feature independence is thought to provide, which was explored in Motivation Subsection 2.3.
@@ 6 Related Work & Future Directions
Alternative approaches to gaussianization, such as transforms for gaussianizing heavy-tailed distributions (Goerg, 2015), iterative gaussianization techniques (Chen & Gopinath, 2000; Laparra et al., 2011), and copula-based gaussianization (Nelsen, 2006), offer interesting directions for future work. Non-parametric techniques for gaussianizing, for example those using quantile functions (Gilchrist, 2000), may not be easily amenable to the deep learning setting where models are trained using backpropagation and gradient descent.
@@ 7 Conclusion
Among the methodological developments that have spurred the advent of deep learning, their success has often been attributed to their effect on the model’s ability to learn and encode representations effectively, whether in the activations or in the weights. This can be seen, for example, by considering the importance of initializing model weights suitably, or by the effect different activation functions have on learning dynamics.
@@
It is perhaps nowhere clearer what representational benefit normality normalization provides, than when considering that no additional learnable parameters, relative to existing normalization layers, were introduced. This highlights – and precisely controls for the effect of – the importance of encouraging models to encode their representations effectively.
@@
This work is of general interest to the machine learning and broader scientific community. There are many potential applications of the work, among which it is difficult to judiciously highlight one such possible application, in place of another.
@@ A.1 Noise Robustness
Table 4 demonstrates the increased robustness to noise obtained when using BatchNormalNorm in comparison to BatchNorm. This substantiates the applicability of the noise robustness framework presented in Motivation Subsection 2.1, and consequently of the benefit of gaussianizing learned representations in normality normalization.
@@ A.2 Speed Benchmarks
The plots show a close correspondence for test-time performance, with a larger deviation at training time. It is worth noting however, that the present work serves as a foundation, both conceptual and methodological, for future works which may continue to leverage the benefits of gaussianizing. We believe improvements to the runtime of normality normalization can be obtained by leveraging approximations to the operations performed in the present form of normality normalization; for example the operations $\log\left(1+h\right)$, and raising to the power.
Figure 11: As in Figure 5 and Figure 6, but for networks at initialization. The plots illustrate graphically that, at initialization, networks with either BatchNormalNorm or BatchNormal have close-to Gaussian pre-activations. However, as the networks are trained, BatchNormalNorm enforces and retains normality while BatchNorm does not, as evidenced by Figure 6.
@@ A.3 Normality at Initialization
Figure 11 shows representative Q–Q plots, together with an aggregate measure of normality across model layers, for post-power transform feature values when using BatchNormalNorm, and post-normalization values when using BatchNorm, for models at initialization. It illustrates graphically that at initialization, the pre-activations are close-to Gaussian regardless of the normalization layer employed; and thus that only the model trained with BatchNormalNorm enforces and maintains normality throughout training, as evidenced by Figure 6. Note that the Q–Q plots presented in Figure 5 and Figure 11 are obtained for the same corresponding minibatch and channel combinations.
@@ Appendix B Lemmas
where the second equality follows because by assumption the marginals are normally distributed, the inequality follows because the normal distribution maximizes entropy, and in the second-last equality we have used the expressions for the entropy of the univariate and bivariate normal distributions. ∎
@@ Appendix C Derivation of the Power Transform Negative Log-Likelihood
As in Section 3, consider a random variable $H$ from which a sample $\bm{h}=\left\{h_{i}\right\}_{i=1}^{N}$ is obtained. Recall that the power transform gaussianizes $\bm{h}$ by applying Equation 3 for each $h_{i}$, where the parameter $\lambda$ is obtained using maximum likelihood estimation (MLE), so that the transformed variable is as normally distributed as possible.
Denote the transformed random variable as $X\sim\mathcal{N}\left(x;\mu,\sigma^{2}\right)$, where $x=\psi\left(h;\lambda\right)$, and $\mu=\mu\left(\lambda\right)$, $\sigma^{2}=\sigma^{2}\left(\lambda\right)$; i.e. $\mu$ and $\sigma^{2}$ are in general functions of $\lambda$. Obtaining the MLE for $\left(\mu,\sigma^{2},\lambda\right)$, shown next, requires evaluating the probability density function (PDF) of $H$, given the transformed variable $X$ is normally distributed.
@@ F.1 ResNet and WideResNet Experiments
We used a variety of residual network (ResNet) (He et al., 2016) and wide residual network (WideResNet) (Zagoruyko & Komodakis, 2016) architectures in our experiments. For all of the experiments except those using the TinyImageNet (TinyIN), Caltech101, and Food101 datasets, models were trained from random initialization for $200$ epochs, with a factor of $10$ reduction in learning rate at each $60$-epoch interval. For the experiments using the TinyImageNet (TinyIN), Caltech101, and Food101 datasets, models were trained from random initialization for $100$ epochs, with a factor of $10$ reduction in learning rate at epochs $40,70,90$. A group size of $32$ was used in all of the relevant group normalization experiments. For the Caltech101 dataset, each run used a random $90$/$10$% allocation to obtain the training and validation splits respectively [^13] The official Caltech101 dataset does not come with its own training/validation split.. Each such run used its own unique random seed to generate the splits for that run, which facilitates greater precision in the reporting of our aggregate results across the runs.
@@ F.2 Vision Transformer Experiments
The training configuration of the model and dataset combinations which use layer normality normalization (LayerNormalNorm/LNN) and layer normalization (LayerNorm/LN) are as follows. We used a vision transformer (Vaswani et al., 2017; Dosovitskiy et al., 2021) model consisting of $8$ transformer layers, $8$ attention heads, hidden dimension size of $768$, and MLP dimension size of $2304$. A patch size of $4$ was used throughout, except in the Food101 and ImageNet experiments, where it was set to $16$.

### v2

@@ 1 Introduction
Our experiments comprehensively demonstrate the general effectiveness of normality normalization, in terms of its generalization performance, its strong performance across various common factors of variation such as model width, depth, and training minibatch size, which furthermore serve to highlight why it is effective, its suitability for usage wherever existing normalization layers are conventionally used, and its effect on improving model robustness under random perturbations.
@@ 2.1.1 Overview of the Framework
Under first and second moment constraints, the normal distribution is at the same time the best possible signal distribution, and the worst possible noise distribution; a result which can be studied in the context of the Gaussian channel (Shannon, 1948), and through the lens of the mutual information game (Cover & Thomas, 2006). In this framework, $X$ and $Z$ denote two independent random variables, representing the input signal and noise, and $Y=X+Z$ is the output. The mutual information between $X$ and $Y$ is denoted by $I\left(X;Y\right)$; $X$ tries to maximize this term, while $Z$ tries to minimize it. Both $X$ and $Z$ can encode their signal using any probability distribution, so that their respective objectives are optimized for.
@@ 2.1.2 Relation to Learning
How might this framework relate to the learning setting? First, previous works have shown that adding noise to the inputs (Bishop, 1995) or to the intermediate activations (Srivastava et al., 2014) of neural networks can be an effective form of regularization, leading to better generalization. Moreover, the mutual information game shows that, among encoding distributions with first and second moment constraints, [^1] Note that conventional normalization layers respect precisely these constraints, since the mean and variance are used to normalize the pre-activations. the normal distribution is maximally robust to random perturbations. Taken together these suggest that encoding activations using the normal distribution is the most effective way of using noise as a regularizer, because a greater degree of regularizing noise in the activations can be tolerated for the same level of corruption.
@@
Finally, there exists a close correspondence between the mutual information between the input and the output of a channel subject to additive Gaussian noise, and the minimum mean-squared error (MMSE) in estimating the input given the output (Guo et al., 2005). This suggests that when Gaussian noise is added to a given layer’s activations, quantifying the attenuation of the noise across the subsequent layers of the network, as measured by the mean-squared error (MSE) relative to the unperturbed activations, provides a measurable proxy for the mutual information between the activations of successive layers in the presence of noise.
@@ 2.3 Maximally Independent Representations
We use these results to motivate the following argument: for a given level of correlation, encouraging normality in the feature representations of units would lead to the desirable property of maximal independence between them; in the setting where increased unit-wise normality also lends itself to increased joint normality.
@@ 3 Background: Power Transform
The parameter $\lambda$ is obtained using maximum likelihood estimation, so that the transformed variable is as normally distributed as possible, by minimizing the following NLL: [^3] To simplify the presentation, we momentarily defer the cases $\lambda=0$ and $\lambda=2$, and outline the NLL for $h\geq 0$ only, as the case for $h<0$ follows closely by symmetry.
@@ 4.2 Additive Gaussian Noise with Scaling
For each input indexed by $i\in\left\{1,\ldots,N\right\}$, during training555 We do not apply additive random noise with scaling at test time. we have $y_{i}=x_{i}+z_{i}\cdot\xi\cdot s$, where $x_{i}$ is the $i$-th input’s post-power transform value, $z_{i}\sim\mathcal{N}\left(0,1\right)$, $\xi\geq 0$ is the noise factor, and $s=\frac{1}{N}\left\lVert\bm{x}-\bar{\bm{x}}\right\rVert_{1}$ represents the zero-centered norm of the post-power transform values, normalized by the sample size $N$.
@@ 5.4 Effectiveness Across Model Configurations
Figure 2 shows that BatchNormalNorm outperforms BatchNorm across varying WideResNet architecture model widths. Of particular note is that BatchNormalNorm shows strong performance even in the regime of relatively small network widths, whereas BatchNorm’s performance deteriorates. This may indicate that for small-width networks, which do not exhibit the Gaussian process limiting approximation attributed to large-width networks (Neal, 1996; Lee et al., 2018; Jacot et al., 2018; Lee et al., 2019), normality normalization provides a correcting effect. This could, for example, be beneficial for hardware-limited deep learning applications.
@@ 5.6 Comparison of Additive Gaussian Noise With Scaling and Gaussian Dropout
The second is additive Gaussian noise, but without scaling by each channel’s minibatch statistics. This corresponds to the proposed method in the case where $s$ is fixed to the mean of a standard half-normal distribution, [^9] This value for $s$ precisely mirrors how it is calculated in Algorithm 1, since recall $s=\frac{1}{N}\sum_{i=1}^{N}\lvert x_{i}-\bar{x}\rvert$. i.e. $s=\sqrt{\frac{2}{\pi}}$ across all channels; and thus does not depend on the channel statistics.
@@ 5.7 Effect of Degree of Gaussianization
Here we consider what effect differing degrees of gaussianization have on model performance, as measured by the proximity of the estimate $\hat{\lambda}$ to the Newton-Raphson solution, which was given by Equation 5.
We control the proximity to the Newton-Raphson solution using a parameter $\alpha\in\left[0,1\right]$ in the following equation
@@
where $\alpha=1$ corresponds to the Newton-Raphson solution, and decreasing values of $\alpha$ reduce the strength of the gaussianization.
@@ 5.9 Additional Experiments & Analysis
Appendix A.1 demonstrates that models trained using normality normalization are more robust to random noise at test time. This substantiates the applicability of the noise robustness framework presented in Motivation Subsection 2.1, and consequently of the benefit of gaussianizing representations.
@@
Appendix A.2 shows that normality normalization increases runtime; with a larger deviation at training time than at test time.
@@
Appendix A.3 provides a graphical illustration of the fact that at initialization, layer pre-activations are close to Gaussian regardless of the normalization layer employed; thus only models trained with normality normalization maintain gaussianity throughout training.
@@
Appendix A.4 demonstrates that normality normalization imbues models with greater joint normality and greater independence between channel features, throughout the layers of a model. This is of value in context of the benefit feature independence is thought to provide, which was explored in Motivation Subsection 2.3.
@@ 6 Related Work & Future Directions
Alternative approaches to gaussianization, such as transformations for gaussianizing heavy-tailed distributions (Goerg, 2015), iterative gaussianization techniques (Chen & Gopinath, 2000; Laparra et al., 2011), and copula-based gaussianization (Nelsen, 2006), offer interesting directions for future work. Non-parametric techniques for gaussianizing, for example those using quantile functions (Gilchrist, 2000), may not be easily amenable to the deep learning setting where models are trained using backpropagation and gradient descent.
@@ 7 Conclusion
Among the methodological developments that have spurred the advent of deep learning, their success has often been attributed to their effect on the model’s ability to learn and encode representations effectively, whether in the activations or in the weights. This can be seen, for example, by considering the importance attributed to initializing model weights suitably, or by the effect different activation functions have on learning dynamics.
@@
It is perhaps nowhere clearer what representational benefit normality normalization provides, than when considering that no additional learnable parameters relative to existing normalization layers were introduced. This highlights – and precisely controls for the effect of – the importance of encouraging models to encode their representations effectively.
@@
Acknowledgments
We acknowledge the support provided by Compute Ontario (computeontario.ca) and the Digital Research Alliance of Canada (alliancecan.ca), and the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), Discovery Grant RGPIN-2021-02527.
@@ Acknowledgments
This work is of general interest to the machine learning and broader scientific community. There are many potential applications of the work, for which endeavoring to judiciously highlight one such possible application, in place of another, would be a precarious undertaking.
@@ A.1 Noise Robustness
Table 4 demonstrates the increased robustness to noise obtained when using BatchNormalNorm in comparison to BatchNorm. This substantiates the applicability of the noise robustness framework presented in Motivation Subsection 2.1, and consequently of the benefit of gaussianizing representations in normality normalization.
@@ A.2 Speed Benchmarks
The plots show that normality normalization increases runtime; with a larger deviation at training time than at test time. It is worth noting however, that the present work serves as a foundation, both conceptual and methodological, for future works which may continue to leverage the benefits of gaussianizing. We believe improvements to the runtime of normality normalization can be obtained by leveraging approximations to the operations performed in the present form of normality normalization; for example the operations $\log\left(1+h\right)$, and raising to the power.
Figure 11: As in Figure 5 and Figure 6, but for networks at initialization. The plots provide a graphical illustration of the fact that at initialization, networks using either BatchNormalNorm or BatchNorm have close-to Gaussian pre-activations. However, as the networks are trained, BatchNormalNorm enforces and maintains normality while BatchNorm does not, as evidenced by Figure 6.
@@ A.3 Normality at Initialization
Figure 11 shows representative Q–Q plots, together with an aggregate measure of normality across model layers, for post-power transform feature values when using BatchNormalNorm, and post-normalization values when using BatchNorm, for models at initialization. It provides a graphical illustration of the fact that at initialization, the pre-activations are close to Gaussian regardless of the normalization layer employed; and thus that only the model trained with BatchNormalNorm enforces and maintains normality throughout training, as evidenced by Figure 6. Note that the Q–Q plots presented in Figure 5 and Figure 11 are obtained for the same corresponding minibatch and channel combinations.
@@ Appendix B Lemmas
where the second equality follows because by assumption the marginals are normally distributed, the inequality follows because the normal distribution maximizes entropy, and in the second-last equality we have used the expressions for the entropies of the univariate and bivariate normal distributions. ∎
@@ Appendix C Derivation of the Power Transform Negative Log-Likelihood
As in Section 3, consider a random variable $H$ from which a sample $\bm{h}=\left\{h_{i}\right\}_{i=1}^{N}$ is obtained. Recall that the power transform gaussianizes $\bm{h}$ by applying Equation 3 for each $h_{i}$, where the parameter $\lambda$ is obtained using maximum likelihood estimation, so that the transformed variable is as normally distributed as possible.
Denote the transformed random variable as $X\sim\mathcal{N}\left(x;\mu,\sigma^{2}\right)$, where $x=\psi\left(h;\lambda\right)$, and $\mu=\mu\left(\lambda\right)$, $\sigma^{2}=\sigma^{2}\left(\lambda\right)$; i.e. $\mu$ and $\sigma^{2}$ are in general functions of $\lambda$. Obtaining the maximum likelihood estimates for $\left(\mu,\sigma^{2},\lambda\right)$, shown next, requires evaluating the probability density function (PDF) of $H$, given the transformed variable $X$ is normally distributed.
@@ F.1 ResNet and WideResNet Experiments
We used a variety of residual network (ResNet) (He et al., 2016) and wide residual network (WideResNet) (Zagoruyko & Komodakis, 2016) architectures in our experiments. For all of the experiments except those using the TinyImageNet (TinyIN), Caltech101, and Food101 datasets, models were trained from random initialization for $200$ epochs, with a factor of $10$ reduction in learning rate at each $60$-epoch interval. For the experiments using the TinyImageNet (TinyIN), Caltech101, and Food101 datasets, models were trained from random initialization for $100$ epochs, with a factor of $10$ reduction in learning rate at epochs $40,70,90$. A group size of $32$ was used in all of the relevant group normalization experiments. For the Caltech101 dataset, each run used a random $90$/$10$% allocation to obtain the training and validation splits respectively. [^14] The official Caltech101 dataset does not come with its own training/validation split. Each such run used its own unique random seed to generate the splits for that run, which facilitates greater precision in the reporting of our aggregate results across the runs.
@@ F.2 Vision Transformer Experiments
The training configuration of the model and dataset combinations which use layer normality normalization (LayerNormalNorm/LNN) and layer normalization (LayerNorm/LN) are as follows. We used a vision transformer (Vaswani et al., 2017; Dosovitskiy et al., 2021) model consisting of $8$ transformer layers, $8$ attention heads, hidden dimension size of $768$, and multi-layer perceptron (MLP) dimension size of $2304$. A patch size of $4$ was used throughout, except in the Food101 and ImageNet experiments, where it was set to $16$.